OpenStudy (anonymous):

find dy/dt where 4x^2-y=100 and dx/dt equals 8 when x=15

OpenStudy (campbell_st):

use the chain run \[\frac{dy}{dt} = \frac{dy}{dx} \times \frac{dx}{dt}\]

OpenStudy (campbell_st):

its related rates and change the equation to \[y = 4x^2 - 100\] so find dy/dx

OpenStudy (akashdeepdeb):

Just differentiate \(4x^2-y=100\) with respect to \(t\). \[\frac{d}{dt}(4x^2-y) = \frac{d}{dt}100\] \[\frac{d}{dt}(4x^2) - \frac{dy}{dt} = 0\] Apply chain rule to the first term: \[\frac{d}{dx}(4x^2) \frac{dx}{dt} - \frac{dy}{dt} = 0\] Thus, \[\frac{dy}{dt} = \frac{d}{dx}(4x^2) \frac{dx}{dt}\] Can you find this now? :)

OpenStudy (anonymous):

so you find the derivative of 4x^2 and plug 15 into x then multiple that answer by 8?

OpenStudy (akashdeepdeb):

Yes.

OpenStudy (anonymous):

the answer should be 960